Here is the question. I even know the answer. However, I do not know how to solve it.
"A woman walks home by traveling 20 meters north, 25 meters east, and then 5 more meters north. If she could walk home along a straight line, how many fewer meters would she need to walk? Round your answer to the nearest hundredth."
The answer is 14.64 meters.

Answer by checkley75(3666)

THE BEST WAY TO SOLVE THIS PROBLEM IS WITH A GRAPH USING X & Y COORDINATES.
START AT THE INTERSECTION (0,0), GO UP (NORTH) 25 UNITS, THEN GO 20 UNITES TO THE RIGHT (EAST) THEN ANOTHER 5 UP (NORTH).
YOU ARE NOW AT (25,25). NOW DRAW A LINE FROM HERE TO THE ORIGIN (0,0) AND YOU HAVE THE HYPOTENUSE OF A 25 BY 25 RIGHT TRIANGLE.
USING THE FORMULA A^2+B^2=C^2 WE HAVE
25^2+25^2=C^2
625+625=C^2
1250=C^2
C=35.36 METERS IS THE LENGTH OF THE WALK HOME.
50-35.36=14.64 METERS IS THE DIFFERENCE BETWEEN THE TWO WALKS.